OpenStudy (anonymous):

Does the point you select matter when your write a point-slope equation? Explain your reasoning using complete sentences.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@DebbieG

OpenStudy (anonymous):

annyone?

OpenStudy (debbieg):

What do you think? Does it matter what point you use? Try a couple of examples. Use a simple line, like say, y=x+1 m=1 Now pick two different points (you can do this by choosing an x, and computing the y). Lets say: (1, 2) and (2, 3) Using the first point, the point-slope form is: y-2=1(x-1) Using the 2nd point, the point slope form is: y-3 = 1(x-2) is the equation for the line any different? if you solve for y, do you have the same equation?

OpenStudy (anonymous):

i solved for sum problems already and my final answer was " yes it does matter because depending on the slope and the points the answer will come out to be different" ????

OpenStudy (debbieg):

Well, show your work for those where you found it to be different, and we can determine where you went wrong. The equation for a line, is the equation for the line. The POINT chosen to cook up the equation does not change the equation. There is only ONE equation for any given line (if it's in y=mx+b form, that is). The equation is unique,, given the slope and y-intercept, so it can't depend on which points you choose to computer it.

OpenStudy (anonymous):

i selected (2,4) and (3,4) and i came up with this answers y=x+1 y=x+1

OpenStudy (anonymous):

thats point slope form and then next question was Does the point you select matter when your write a point-slope equation? Explain your reasoning using complete sentences. so im guessing it doesnt matter what point you select the answer would be the same?

OpenStudy (debbieg):

Now hold on a second. You "selected" (2,4) and (3,4) how? what is the equation for the line that you are using? Those two points are NOT on the line y=x+1, only the SECOND one is.

OpenStudy (debbieg):

If you are using my example, then (2,4) is NOT on that line. If x=2, y=3 when the equation is y=x+1

OpenStudy (anonymous):

nono im using my example. my points are (2,3) and (3,4) and the quations are y-4+1(x-3) y-3=1(x-2)

OpenStudy (anonymous):

after that i came up with this answers y=x+1 y=x+1

OpenStudy (anonymous):

i came up wit another answer to the actual question "no, the point doesnt matter when youre writing a point slope equation as long as the point lay in the same line" ???

OpenStudy (debbieg):

Yes, that's correct. Those two points (not the ones you originally stated) are both on the line y=x+1. As you can see, regardless of the point you use in the formula, the equation for the line will be the same.

OpenStudy (anonymous):

so my answer is correct right!? "no, the point doesnt matter when youre writing a point slope equation as long as the point lay in the same line" thank you!

OpenStudy (anonymous):

@DebbieG

OpenStudy (debbieg):

Yes, that's correct. As I said above. :)

OpenStudy (anonymous):

thank you!

OpenStudy (debbieg):

you're welcome :)

OpenStudy (anonymous):

last question can you please hem me with this one? Explain, using complete sentences, why it is important to understand any limitations on the domain and range and sorry i was shoering